OpenStudy (oleg3321):

Given the function g(x) = 8x - 2, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values.

OpenStudy (samsan9):

don't you just plug in the numbers and then just choose which one is the correct ones?

OpenStudy (triciaal):

did you find g(-2) and g(4) then look at the choices to see what option is true?

OpenStudy (triciaal):

g(x) = 8x - 2, g(-2) = 8(-2)-2 = -16 -2 =-18 g(4) = 8(4)-2 = 32-2 = 30 g(-2) is smaller than g(4)

OpenStudy (triciaal):

yes

OpenStudy (triciaal):

(x, y) (3n - 4, f(3n - 4))

OpenStudy (oleg3321):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function?

OpenStudy (oleg3321):

-2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180,

OpenStudy (oleg3321):

@triciaal

OpenStudy (oleg3321):

@JMark

OpenStudy (oleg3321):

@aum

OpenStudy (oleg3321):

@doulikepiecauseidont

OpenStudy (anonymous):

Post a new thread